What is the inductance of an LC circuit with  

C = 3.30 µF

oscillating at

59.0 Hz?


H
Answer #1

the inductance of the inductor is,

L = 1/ 4*pi2*f2*C = 1/ 4*pi2*59*59*3.3X10-6 = 2.2 H